A company manufactures a new type of mosquito repellents with the period of effectiveness that is approximately normally distributed with a standard deviation of 30 hours. A sample of 40 repellents was taken and the average period of effectiveness is 290 hours. Calculate a 96% confidence interval for the true mean of the period of effectiveness


1
Expert's answer
2022-06-02T09:06:46-0400

The critical value for "\alpha = 0.04" is "z_c = z_{1-\alpha/2} = 2.0537."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(280.258, 299.742)"

Therefore, based on the data provided, the 96% confidence interval for the population mean is "280.258 < \mu < 299.742," which indicates that we are 96% confident that the true population mean "\mu" is contained by the interval "(280.258, 299.742)."